OpenStudy (anonymous):

I am lost.... please help! Rational Expressions are kicking my tush! The problem is: 1/x+1/3=1/2

OpenStudy (bahrom7893):

okay workin on it

OpenStudy (bahrom7893):

First multuply everything through by x.

OpenStudy (bahrom7893):

*multiply: 1 + x/3 = x/2

OpenStudy (bahrom7893):

Now multiply everything by 3: 3 + x = (3x)/2

OpenStudy (bahrom7893):

Now multiply everything by 2: 6 + 2x = 3x

OpenStudy (bahrom7893):

*it would be easier to just multiply everything by 6, but I'm just showing all in-between steps.

OpenStudy (bahrom7893):

After: 6 + 2x = 3x, subtract 2x from both sides: 6 + 2x - 2x = 3x - 2x

OpenStudy (bahrom7893):

6 = x; or x = 6 DONE!

OpenStudy (anonymous):

Thank you so much! This class is kicking my tush and I am not sure I am going to make it through it!

OpenStudy (bahrom7893):

ur welcom!

OpenStudy (bahrom7893):

*e